OpenStudy (mackenzie2013):

Out of 200 apples in a basket, 5 are rotten. What is the experimental probability that an apple chosen at random is a good one? If there 800 apples in a basket, how many of them are likely to be good? 95%; 720 apples 97.5%; 780 apples 2.5%; 20 apples 80%; 615 apples

OpenStudy (mathmale):

@mikey: Please DO NOT HAND OUT ANSWERS! If you want to be of help, then provide Mackenzie with just enough help so that he or she can solve this problem himself / herself.

OpenStudy (mackenzie2013):

I just need my answers checked lol. I chose B.

OpenStudy (mikey):

200 apples and 5 are rotten so 800/200=4 5*4 yes you are correct

OpenStudy (mikey):

I gave u the answer and mathmale told me to not tll u the answer

OpenStudy (mathmale):

"800/200=4 5*4" is not correct. What percentage of the 200 apples are good? Mikey: Again, do NOT give out answers; if you want to help, guide Mackenzie in the problem solving process so that he or she can find his or her own answer. It's particularly serious if you provide direct answers and those answers turn out to be wrong.

OpenStudy (mathmale):

What is the experimental probability that an apple chosen at random is good? What is the experimental probability that an apple chosen at random is bad? Show your work, please.

OpenStudy (mackenzie2013):

He was correct.

OpenStudy (mackenzie2013):

I needed an answer checked, I didnt see it as giving out an answer. :)

OpenStudy (mathmale):

I disagree. It's very unwise to accept someone else's work as your own. What does "800/200=4 5*4" represent, if not an answer? What is the experimental probability that an apple chosen at random is good? What is the experimental probability that an apple chosen at random is bad? Show your work, please.